OpenStudy (purplemexican):

If f(x) = 1 – x, which value is equivalent to |f(i)|?

OpenStudy (anonymous):

f(i) = 1 - i For the complex number \(z=a+bi\), \[|z|=\sqrt{a^2+b^2}\] \[|1-i|=\sqrt{(1)^2+(-1)^2}\]